Report of Activities on the RVIB N.B. Palmer Cruise 02-04

7 August 2002


August 7 was our last day in Crystal Sound.  In the wee hours, a CTD transect was run along the southern portion of the sound out to the entrance of Matha Strait.  At the final CTD station location, a combined ice collection operation and an ROV under-ice survey were done.  The Palmer then steamed back into Crystal Sound toward the location where the Tucker Trawl was done yesterday.  The 1-m MOCNESS was accomplished successfully, but not without difficulty.  The pressure sensor froze up during the extended setup time on the deck typical of a first tow and did not thaw out until about 20 minutes into the tow.  On one occasion, the towing wire was caught by a large ice chunk that drifted into the wake and the ship had to stop and back up to free the cable. The same thing occurred later in the evening when the 10-m MOCNESS was towed for the first time on this cruise, but it did not affect the ultimate success of either tow. Both net systems had very nice catches in their respective nets, which showed a complicated vertical structure of the plankton and nekton living in the water column. The significant concentrations of krill in the Crystal Sound area at this time of year is likely a major reason why the many marine predators were present.  We completed the work in Crystal sound in the late evening and then headed out Matha Strait to the west with the L.M. Gould close behind. 


In general, the work in Crystal Sound was quite successful. All of the instruments were deployed for the first time, and with only a couple of exceptions worked, just fine.  The microstructure instrumentation has not yet worked when deployed on the CTD, but the trouble-shooting has turned up several fixable problems and the system should be operational when next put into the water.  BIOMAPER-II suffered a failure of the transmitting circuitry for the 420 and 1000 kHz frequencies during the period of calibration two days ago. Efforts to determine what component failed have not yet been successful, but are ongoing. The lower three frequencies remain operational and the system should be survey-ready at the start of the grid work.


The benign weather pattern we have experienced over the past couple of days, dominated by a high pressure system, continued on 7 August.  Although the air was cold — mostly between -8 and -10ºC — the winds were light out of the southwest, so work on deck was comfortable. The barometer continued to fall slowly and varied from 1015 mb in the early morning to 1010 mb in the evening. Sea surface temperature remained at the freezing point (-1.845ºC)


CTD Group report (Baris Salihoglu, Eileen Hofmann, Bob Beardsley, Chris MacKay,  Francisco (Chico) Viddi, Sue Beardsley)

In the morning of 7 August, we completed a small-scale CTD survey in Crystal Sound that included 7 CTD stations spaced along two transects. The first transect, which extended North to South across Matha Strait, was finished in the morning of 6 August.  We waited until the late evening of 7 August to start the second transect across Crystal Sound because the daylight hours were reserved for studies that require light, such as bird and cetacean observations. The CTD group was also busy during the daytime. Four members of the group (Eileen Hofmann, Bob Beardsley, Sue Beardsley, and Baris Salihoglu) joined the Sea Bird group in the afternoon of 6 August to help them diet-sample some Adélie penguins. Thanks to Chris Ribic and Eric Chapman for letting the physical oceanographers enjoy some predator work.


The second transect across Crystal Sound, consisted of 4 stations, the deepest of which were the two central stations. The CTD casts were made to the bottom at all stations. The CTD data from the first three casts had to be recovered from the back-up tapes because the data acquisition computer had been corrupted by a virus.


Vertical temperature distributions constructed for both transects, showed maximum values of 1.3ºC and 1.4ºC at the bottom of the deepest stations (590 m and 860 m). Temperatures below ~250 m were above 1.0ºC indicating that the trough that extends into Crystal Sound from the continental shelf was filled with Circumpolar Deep Water (CDW) that has been modified by mixing with the overlying Antarctic Surface Water.  The Winter Water layer extended over the upper 80 m to 100 m along both transects. Surface water temperatures were slightly above freezing at about -1.75°C along both transects except for the Northern part of the first transect (towards Watkins Island) where temperatures of -1.8°C and below were observed.


The salinity values reached 34.72 at the deepest points along both transects, which indicated the presence of Upper Circumpolar Deep Water. The vertical salinity distribution along the second transect showed that the isohalines above 250 m were tilted upwards towards the west, indicating a possible flow into Crystal Sound through Matha Strait. This is consistent with concurrent ADCP observations.  The potential density profiles for both transects were consistent with the salinity profiles indicating that salinity is the dominant factor determining density in this region.  This is typical for Antarctic waters.


Sea Birds (Chris Ribic and Erik Chapman)

The ship worked an area about 10 to 15 miles south of the Barcroft Islands today (7 August), where the ice was consistently around 9/10ths coverage.  The larger leads that were found near the Barcroft Islands were absent from this area and no Adélie Penguins were observed today.  Flying birds were also virtually absent, and just a small number of Snow Petrels were observed from the bridge today.  It appeared that within Crystal Sound, Adélie Penguins were more abundant in the lighter ice cover near the Barcroft Islands.


Like the breeding colonies near Anvers Island and Adelaide Island, the colonies in the Barcroft Islands are near a large, glacially scoured trough that is at least 1,000 meters deep.  It is believed that these troughs are important for krill reproductive biology, and may explain the location of Adélie breeding colonies.  Presumably by breeding near these troughs, Adélies have access to large concentrations of krill during a period of high energy demand during breeding in the austral spring and summer.  A 10-meter MOCNESS tow in this trough revealed a large amount of zooplankton, including Antarctic krill (Euphausia superba).  This is the species of krill found in the penguin diets taken yesterday and found at depths within the diving range of Adélies.  These results suggest that these troughs may be areas of high krill density in the winter, as well.


Marine Mammal report (Chico Viddi)

Weather conditions the last two days (7 & 8 August) have not provided much opportunity to make marine mammal observations. During this period, the ice ranged from 9 to 10/10ths coverage and the days have been dark, dreary, and cold. Poor visibility has been the key factor in limiting observation hours, which only totaled 4.67 hr.  Nevertheless, 7.83 “incidental” hours of observation were made. The cumulative hours of marine mammal observations to date now total 26 (excluding incidental hours). The most frequent marine mammal seen on Tuesday, 7 August, was the crabeater seal (Lobodon carcinophagus).  No cetaceans were observed, but a fantastic pinniped sighting was made during the penguin diet-sampling (-66º 28.168′S; -66º 01.057′W). Once the three penguins were caught, the rest of the birds (seven of them) escaped further away and stayed inactive while recovering from “the human attack”. Unfortunately for them, a leopard seal (Hydrurga leptonix) surfaced right next to the ice floe they were on. The penguin reaction towards the leopard seal's presence was evident. Leopard seals are known to be facile hunters of penguin and seal pups.  They hide underwater below the ice and wait until penguins go into the water. Little is known about leopard's attacks on ice.


Many crabeater seals were seen on 7 August with the first sighting at 1130, 9º to starboard and 1100 m from the vessel (-66º 32.44′S; -67º 30.312′W).  About 18-20 seals were counted on ice floes in an area of about 1 km2. A few more seal sightings were made during a transit to a starting point for a 1-m MOCNESS tow. During the tow, a minke whale (Balaenoptera acutorostrata) was seen at 1605, 36º to port and 870 m from the vessel (-66º 31.66′S; -67º 11.27′W, first sighting was made by Erik Chapman). The whale swam slowly for a couple of minutes among the ice floes, in very small areas of open water. About 12 crabeater seals were on the ice close to where the whale was and at least five seals were in the water less than 10 meters from the whale.


Microplankton report (Phil Alatalo, Gusatvo Thompson, Dicky Allison, and Scott Gallager)

The objectives of our microplankton studies are: 1) to provide an additional perspective on the microplankton prey field utilized by larval and adult krill, by quantifying abundance and motion characteristics, (i.e., swimming behavior) in relation to particle size distribution; 2) to determine the vertical and horizontal distribution of autotrophic and heterotrophic microplankton, including pelagic ciliates, heterotrophic dinoflagellates, and silicaflagellates along the western Antarctic Peninsula during austral autumn and winter; 3) to relate microplankton distributions to vertical gradients in density, salt, mixing intensity, and light distribution, and horizontal gradients in water mass distribution and surface currents; and 4) in collaboration with Kendra Daly’s group, to determine experimentally the rates of larval krill feeding on microplankton and detritus.


We are using a variety of sampling techniques in this study. First, we use the standard method of collecting water with a 10-liter Niskin bottle which will allow us the quantify microplankton along the CTD station grid extending about 20 nm both north and south of Marguerite Bay and 20 nm offshore.  CTD bottle depths for microplankton sampling are chosen keeping the following vertical regions of the water column in mind: the upper mixed layer, a fresher water lens (if present usually <20 m), the halocline beneath the mixed layer, chlorophyll maxima and minima, and usually a near-bottom deep sample.  Four samples are taken at each CTD station while more are taken if specific regions or strata seem interesting based on data from the CTD, BIOMAPER-II or the VPR.  Samples are removed from the top of the Niskin bottles by gently siphoning through wide bore tubing. We have shown that this procedure minimizes damage during sample transfer particularly to large protests and aggregates (marine snow). Each sample depth is processed by preserving 400 ml in 2% acid Lugol's fixative and by observing swimming behavior on live, unconcentrated samples in a specially developed optical  recording device.  For the purpose of distinguishing between heterotrophs and autotrophs, 200 to 1000 liter samples are processed by filtration onto 0.45 μm black polycarbonate filters and stained with DAPI, FITC or acridine orange.  Slides are held at 0ºC in the dark for a few hours until observed under fluorescence microscopy using a DAPI filter set on a Zeiss Axiophot upright microscope with 20x and 40x objectives. Digital images are saved for further counting and processing.  Using DAPI stain, large heterotrophic protists appear blue with white nuclei, while diatoms, autotrophic dinoflagellates, autotrophic, mixotrophic, and other pigment-containing cells appear also to contain a low level of orange or red fluorescence. Automated particle tracking of microplankton from video recordings in the optical system requires capturing a 30 s video sequence at 30 frames per second into an AVI file, followed by importing the AVI into a Matlab program one frame at a time. Each frame is binerized against a threshold, and each “in-focus” particle's centroid and maximum and minimum axes are recorded in a matrix. The next frame is imported and a second matrix of pixel locations is produced.  A simple nearest-neighbor algorithm is then used to determine if there are particles within a certain displacement window between matrix one and matrix two.  If the centroids are within the window, a particle path is created.  After all paths have been created the ensemble mean velocity vector for all particles in each frame is subtracted from the instantaneous velocity vector of each particle in the field.  This process removes any common mode movement associated with ship roll.  The result of the processing is a table of data for each particle in the field for calibrated diameter, displacement, speed, motion vector, net to gross displacement (NGDR), and  energy dissipation (calculated  by the Lagrangian integral length scale technique).  These statistics are used as characteristics in a discriminant analysis to determine associations between the swimming behavior of microplankton.  The result is a description of the prey field from the perspective of the energy and frequency of motion and size distribution.


Feeding studies of larval krill foraging on microplankton will begin as soon as the divers are able to collect significant quantities, hopefully within the next day. Observations from the video recording system and selected microscope slides show a dramatic change from the previous GLOBEC-III cruise.


Virtually all large diatoms, dinoflagellates, and protozoans are absent from the waters of Crystal Sound and Matha Strait.  Very small particles predominate all sampled depths of the water column.  A few small flagellates have exhibited most of the recorded motility.  Microscopic examination shows some small diatoms (Rhizosolenia sp., Fragilariopsis sp., Chaetoceras sp.), particularly at depth, and a remarkable amount of fluorescent debris, often fibrous, in the surface waters.  Data calculated from the motion analysis confirms the presence of very small particles (15 μm diameter) exhibiting little distinct motion.  The CTD trace from one sample in Matha Strait and seven in Crystal Sound show quite similar water structure with chlorophyll levels at very low levels throughout the cast.  A sub-surface sample collected at the ice station in Crystal Sound on 6 August was similar in composition and motility to samples collected from the CTD rosette.


MOCNESS Report (Phil Alatalo, Peter Wiebe, Dicky Allison, Ryan Dorland, Scott Gallager, Gareth Lawson)

The 1-m2 MOCNESS (Multiple Opening/Closing Net and Environmental Sensing System) is used onboard to collect zooplankton at discrete depths, while acquiring physical and biological information from the surrounding water.  An Optical Plankton Recorder was added to the standard package for recording size distributions of particles as they flow past the instrument.  To enhance the capture of strong swimming krill, a high-powered underwater strobe light was also mounted on the MOCNESS frame to temporarily blind zooplankton with good visual capability.  Nine nets collect zooplankton for genetic analysis, population structure, species identification, and quantification of abundance and biomass for ground-truthing acoustic sensors used on the ship.  All samples are preserved onboard for further analysis in the laboratory at home.


The first MOCNESS tow occurred on 7 August in Crystal Sound at approximately 1400.  Despite a temporarily frozen pressure sensor that thawed out after about 20 minutes into the tow and began to provide good depth information, and a few incidents where the towing cable was snagged and carried backwards by ice chunks drifting into the wake region, all systems eventually functioned correctly and the tow was successfully conducted.  Krill (all stages) were abundant and ctenophores present at depths of 200-300 m, while copepods, chaetognaths, and amphipods became significant at depths of 150 m to 50 m.  Krill were present up to 50 m, and virtually nothing was caught above 25 m.  Many of the observed krill were Thysanoessa sp.


Current position and conditions

We have nearly completed the first day of our steam towards the southern sector of the SO GLOBEC grid with the L.M Gould following in our wake. Our current position at 2327 on 8 August is -66º 41.839′S; -68º 56.651′W. The air temperature is -6.4ºC and the sea temperature is -1.839ºC. Both anemometers were serviced by Todd Johnson who climbed the science mast and removed the thick coating of snow and ice from the anemometer propeller blades and they are again working. Winds are out of the northeast (055) at 20 to 25 kts. The barometer is at 999.7 and falling with the approach of a low pressure system. The skies are cloudy and a light snow is falling.


Cheers, Peter